# A Review of Community-Centric Power Systems Resilience Assessment and Enhancement Strategies

### v1

@@ 3 Resilience of Interdependent Power and Community Infrastructure Systems
Modern power systems operate within an interconnected ecosystem of critical infrastructures whose joint performance determines overall resilience. While power system resilience traditionally refers to the grid’s ability to withstand and rapidly recover from high-impact, low-probability (HILP) events while maintaining operational stability and infrastructure integrity [76], community functionality and broader infrastructure recovery are inherently linked to this capability. Community resilience—defined as the capacity of populations and institutions to prepare for, absorb, and recover from disasters—depends heavily on the continuity of electricity supply and the robustness of interdependent systems such as transportation, water, gas, and telecommunication networks [75]. In highly electrified and infrastructure-reliant societies, disruptions in the power grid can cascade across sectors, amplifying socioeconomic and operational impacts. Accordingly, assessing resilience now requires a system-of-systems (SoS) perspective that captures both technical inter-dependencies and community consequences. Table 1 summarizes representative studies examining the interdependence among the electric power systems and other communities and critical infrastructures.
3.1 Socioeconomic and Behavioral Dimensions
A growing body of research has quantified how socioeconomic and behavioral factors shape outage exposure, duration, and subsequent recovery. A data-driven Social Vulnerability Index (SoVI) framework in [70] combined Principal Component Analysis (PCA), Adaptive Random Forests (ARF), and variance-based sensitivity analysis (VBSA) to identify the dominant societal drivers of outage vulnerability. Projected SoVI values for 2030 indicate a reduction in vulnerability disparities compared to 2022, reflecting long-term adaptation and incremental infrastructure improvement. Similar disparities were documented in [113], where a 0.1 increase in socioeconomic vulnerability was associated with a 6.1% increase in outage duration. Counties with the longest restoration delays (+6 hours above the mean) exhibited very high SoVI values ($\text{SoVI}\approx 0.91$).
Outage exposure also differs across racial and ethnic lines. A multilevel Bayesian analysis in [114] showed that counties with high racial/ethnic vulnerability $(\text{SoVI}>0.9$) experienced the longest outage durations during extreme weather events, including hurricanes, heat waves, and floods. Similarly, a spatially resolved analysis in [115] found that counties with high SoVI experienced approximately two times more outages lasting $\geq 1$ hour and approximately three times more outages lasting $\geq 8$ hours compared to low-SoVI counties.
Behavioral factors further amplify these disparities. A survey study in New York City [72] found that although 58% of respondents reported feeling generally prepared for outages, only 32% felt capable of enduring multi-day interruptions. Users of durable medical equipment (DME) were more prepared (70%), yet 74% still feared serious health complications during outages. These findings reveal a crucial gap between perceived and actual preparedness, especially for medically vulnerable or socially isolated residents. Additional household-level evidence from [116] shows that minority households reported significantly longer outages, while lower-income households experienced almost twice as many outage events. Table 1 summarizes the interdependency between power systems resilience and SoVI.
Age and health status further influence outage-related risk. Children, older adults, and medically fragile populations experience disproportionately severe impacts from prolonged outages and associated cascading failures [73, 71]. Documented public health risks include cardiovascular events, respiratory complications, renal failure, heat stress, gastrointestinal illness, and increased mental health burdens as outage durations rise [73]. To mitigate these outcomes, recent work highlights the importance of community-scale renewable energy resources and medically focused resilience planning to maintain essential electricity access during HILP events [71]. At the same time, emerging socio-technical models increasingly incorporate behavioral and human-response data, enabling AI-driven forecasting of consumer actions under uncertainty [79].
@@ 3.2 Infrastructure Interdependence and System-of-Systems Perspective
Modern power systems are part of a multilayer infrastructure network in which failures can cascade rapidly from one subsystem to another [101]. As illustrated in Figure 5, the electric grid is tightly coupled with telecommunications, natural gas and heating networks, banking and finance, transportation, water systems, emergency services, and government operations [77, 78]. Each subsystem retains partial operational independence yet depends on shared flows of energy, information, and logistics for sustained functionality. Under normal conditions, such coupling enhances system efficiency through coordinated control, information sharing, and optimized resource allocation. However, during HILP events, these same inter-dependencies accelerate disruption propagation and amplify cascading failures across sectors [80].
@@
From a resilience perspective, these coupled infrastructures represent both stabilizing supports and potential pathways for systemic collapse. Telecommunications enable situational awareness and remote control; gas and water networks supply essential fuels and cooling resources for generation; and transportation systems facilitate repair crew mobilization, logistics, and power system recovery activities [100]. Consequently, resilience evaluation must adopt a SoS viewpoint that captures interdependence across sectors instead of assessing the power grid in isolation. Recent studies highlight this need by quantifying cross-sectoral sensitivity, recovery delays, and performance degradation under multiple hazards.
@@
Water infrastructure exhibits strong dependence on electricity for pumping and pressure maintenance. Optimization-based resilience modeling in [117] showed that placing distributed generators (DGs) near pump stations can increase Water Distribution System (WDS) resilience by 13.35%, while local backup generation yields an additional 3.58% improvement. Under more severe disruptions, potable water systems experience drastic degradation when electrical service is lost. Using the Water Network Tool for Resilience (WNTR) with Pressure-Dependent Demand modeling, [119] demonstrated that hurricane-induced power outages reduce potable water system resilience by nearly 50%. These findings underline the critical reliance of water infrastructure on grid stability and the importance of backup power for maintaining essential public health services.
@@
Interdependency between electricity, natural gas, and heating systems is equally impactful under extreme events. Dynamic optimal energy flow modeling in [118] showed that incorporating gas storage improves gas system resilience by 29.5% and that adding backup generators increases heating-system resilience by 15.6% following hurricanes. Multi-hazard Monte Carlo simulations in [120] further revealed that wind and earthquake events reduce outage energy resilience from 89% (wind) and 83% (earthquake) to 75%, accompanied by a 76%.
@@
Transportation infrastructure also relies heavily on the power grid for operational continuity. Seismic resilience simulations in [75] showed that increasing earthquake magnitude from $M=6$ to $M=8$ reduces transportation network resilience by nearly 50%. Complementary robust optimization studies in [74] demonstrate that strategic hardening of critical power lines and placement of DGs significantly reduce disruption costs and preserve the operability of transportation-critical loads such as traffic signals, particularly during HILP events.
@@
Linear regression analysis of post-hurricane recovery in [121] quantified strong cross-sector dependencies during restoration. Water, telecommunications, and hospital systems relied on power recovery for 89%, 93%, and 77% of their restoration trajectories, respectively. These findings affirm that power restoration serves as the primary driver of cross-infrastructure recovery and underscore the dominant role of electricity as the backbone of community resilience.
@@ 3.3 Resilience Hubs for Community Infrastructure Support
Resilience hubs provide a comprehensive, community-centered approach to addressing these multifaceted vulnerabilities. They recognize the interconnected nature of critical community infrastructure systems and the socioeconomic pathways through which HILP events affect community members. Resilience hubs also integrate public health and safety goals with technical considerations for power, transportation, communications, and essential services.
Resilience hubs are community-oriented facilities, such as schools, community centers, or libraries, that provide everyday public services during normal operating conditions and serve as disaster shelters, resource centers, and infrastructure support during emergencies (Fig. 6). During extreme events, they provide protected environments with physical safety, reliable power, thermal comfort, medical support, communications access, and other essential services [122].
@@
Beyond these community-facing roles, resilience hubs reflect the interdependence of community and grid resilience, addressing both immediate individual social vulnerability concerns and technical system-level performance issues [123]. A resilience hub may focus on supporting medically vulnerable populations, stabilizing public health services, ensuring access to cooling or clean water, or enhancing grid reliability using Distributed Energy Resource (DER) and adaptive load management. In doing so, resilience hubs directly address the intertwined technical and social factors that shape disaster impacts and recovery.
Resilience hubs are equipped with resilient backup power systems that not only ensure uninterrupted on-site access to critical services but can also provide grid-supporting functions [123]. This flexible capacity can stabilize nearby interdependent community infrastructure by supporting water pumping stations, emergency communications, transportation assets, and reliable electric vehicle (EV) charging during outages or evacuations. As a result, resilience hubs can be strategically located either to optimize accessibility for residents [124] or near critical infrastructure such as transportation corridors, water distribution systems, gas networks, and emergency services to help minimize cascading failures during grid disruptions [125, 123].
The framework in [124] designs and evaluates energy systems for resilience hubs in a region prone to extreme heat, identifying communities with limited access to air conditioning and heightened vulnerabilities to power outages. The resulting resilience hub design incorporates a DER portfolio sized to maintain safe indoor temperatures during a heatwave for up to four days without relying on the grid for power. Similarly, the grid-aware optimization in [126] demonstrates how hub-centered microgrids can mitigate outage impacts by balancing local resilience benefits with broader system constraints, demonstrating the role of hubs as both community assets and grid-supporting resources. These examples illustrate how resilience hubs bridge the gap between technical grid robustness and community resilience.
@@
4 Power Systems Resilience Enhancement Strategies
Resilience assessment provides quantitative methods to identify vulnerabilities and strengths of power systems, enabling proactive strategies to strengthen resilience before extreme events occur. Here, different types of resilience-enhancing strategies are identified and explained.
4.1 Hardening Power Network
System hardening refers to modifications made to power system infrastructure to minimize vulnerability to severe events [16, 12]. Hardening is particularly effective at mitigating the impact of extreme events in distribution systems. Typical hardening techniques include reinforcing substations, upgrading or undergrounding power lines, and armoring poles [16, 12, 30]. The U.S. National Electric Safety Code requires the replacement of poles when their structural ability falls below two-thirds of their original strength [30].
In [18, 39], a tri-level optimization is proposed that focuses on hardening power networks to enhance the resilience of distribution systems. It explores strategic decision-making for system hardening through scenario planning under maximum-damage conditions. System hardening, while serving as an effective resilience enhancement strategy, involves a significant financial burden [16, 12]. Consequently, substantial research has focused on developing cost-effective hardening solutions without compromising safety or efficiency [12]. The integration of DER into modern power systems introduces additional challenges. The uncertainty associated with renewable energy operation complicates traditional scenario-based hardening techniques, as it increases the number and severity of system contingencies and potential worst-case scenarios [30].
4.2 Resource Allocation
Resource allocation is a resilience enhancement approach that involves strategically deploying DG units, mobile power sources, and other critical resources at key nodes in the power system to bolster resilience [42]. Given the uncertainty of DER operation during extreme weather conditions, it is essential to back up the lost power to maintain system stability. This helps strengthen grid resilience while maintaining reliable operation during both normal and extreme conditions [40]. Strategically positioning DERs within the network can reduce the need for critical load curtailment and enhance overall system resilience [42, 43]. In [41], a two-stage stochastic mixed-integer programming model was developed to optimize the location and sizing of DGs (in the first stage), minimize the operating costs of DGs and load shedding (in the second stage). Other design decisions, like optimal DG size, can be made using similar optimizations with the objective to maintain resilience above a specified threshold [27].
4.3 Optimal Scheduling of Repair Crews
After an extreme incident, quick repairs of the power system are needed to minimize load-shedding losses. Maintaining resilience requires a proactive repair crew deployment approach to prepare for post-disaster repair and restoration. Optimization models have been proposed to minimize load shedding by sending repair crews to the ideal locations before extreme events. [44]. Their goal is to maximize the expected restored load while minimizing the repair duration. In [45], Anmar Arif et al. propose co-optimization of repairs, reconfiguration, and DG dispatch in a two-stage method for outage management.
4.4 Network Reconfiguration
Network reconfiguration refers to the modification of the network topology by opening normally closed sectionalizing switches and closing normally open tie switches. In order to follow a set of operation constraints, network reconfiguration problems are typically modeled using various objectives, like minimum switching operations, operation costs, or power loss, and are subject to load balancing and network constraints [28]. A number of technologies are employed to improve reconfiguration capacity and distribution system resilience, including smart transformers, on-site DG units, integration of DERs, fault prevention devices, and fault detection systems [12]. Network reconfiguration can be particularly effective at improving power system resilience when more sophisticated measuring instruments are used. Devices like distributed remote controllers, phasor measurement units (PMUs), communicating relays, reclosers, advanced supervisory control and data acquisition (SCADA), and outage management systems (OMS) allow for remote reconfiguration, which can reduce load interruption times. A more effective approach uses the real-time status of the system during each stage of an event to guide subsequent decisions [30].
5 AI Methods for Power Systems Resilience
AI and ML methods are emerging as powerful power system operation and planning tools applied to a variety of tasks such as load forecasting, outage prediction, renewable generation prediction, demand response management [47], anomaly detection [89], stability assessment, control, and restoration [46]. Most ML techniques can be categorized into four groups: 1) Traditional ML, 2) Artificial Neural Network (ANN), 3) Reinforcement Learning (RL), and 4) Advanced Generative and Language Models. Table 2 summarizes different machine learning approaches for power system resilience, highlighting their performance, limitations, and related extreme events.
Power outage prediction has emerged as a critical area of research to bolster the resilience of energy systems against various disruptions, including weather events, equipment malfunctions, and grid instabilities. Logistic regression is a simple, fast, and robust method for outage prediction [51].
Some techniques offer higher accuracy than logistic regression. For instance, Support Vector Machines (SVM) has been utilized to predict outages by assessing the condition of grid components, with SVM generally outperforming logistic regression in terms of accuracy [50, 51]. In [50], a decentralized framework is proposed to isolate affected regions during hurricanes and wildfires. One other notable method is to use ensemble learning techniques. For example, ADABOOST+ has been used to analyze the impacts of lightning and wind on power outages across four cities in Kansas [49]. ADABOOST+ has outperformed ADABOOST.RT, traditional regressional models [86, 51], and neural networks [85].
Convolutional Neural Networks (CNN) combined with long-short-term memory (LSTM) have been applied to predict outages across selected counties in New York, studying the socioeconomic impacts on vulnerable populations based on household income and racial demographics [52].
In [48], the authors enhance prediction model accuracy by pre-classifying events into three severity levels using Quantile Weight Distance (QWD): low (<< 100 outages), moderate (100–1,000 outages), and high (>> 1,000 outages). A conditioned outage prediction model (OPM) is then developed using a Bayesian Additive Regression Tree (BART), Random Forest (RF), and Ensemble Decision Tree (ENS) to outperforms the standard OPM. Addressing data preprocessing challenges, a comprehensive approach for data collection and feature selection is proposed, which includes environmental features, infrastructure specifics, drought, soil moisture, pressure, temperature–humidity, precipitation, convection, and wind [53]. The study suggests that the combination of log-transform, synthetic minority over-sampling technique (SMOTE), and conditional bias correction (ModelF) provides higher accuracy (above 80%). The role of socioeconomic variables in power outage prediction is also examined [88].
Category Methodology Key Findings Drawbacks Extreme Event
Advanced Generative and Language Models Transformer-based [99, 96] Reduce forecast errors compared with numerical models. Event-based (outage count): 84-99% [99]. Self-attention mechanism is limited Natural disaster, cyber-attack.
LLM-based Models [97, 98, 96] Extends transformer-based model, achieve 90–99% accuracy through large-scale parameter tuning and contextual adaptability. Event-based (customer affected): (>>90%) [97, 98]. High computational and storage cost Hurricanes, cyber-attacks, natural disasters.
RL has emerged as a prominent technique in stability control and restoration tasks. In RL applications, the environment is typically modeled as a finite-horizon Markov decision process, defined by the tuple $(S,A,p,r,d_{0},\gamma)$. Here, $S$ and $A$ represent the state and action spaces, respectively. $p(s^{\prime}\mid s,a)$ is the state transition function indicating the probability of transitioning to state $s^{\prime}$ from state $s$ after taking action $a$; $r(s,a)$ is the reward function for taking action $a$ in state $s$; $\gamma$ is the discount factor for future rewards, and $d_{0}$ is the initial state distribution. The results demonstrate that deep reinforcement learning (DRL) outperforms the greedy algorithm and can control voltage regulation [54].
Different RL tasks are characterized by unique state and action spaces, along with task-specific reward functions. [55, 56, 57] focused on using RL for MG formation post-disaster, whereas [58, 59] further explore MG operation. These models often have discrete action spaces (e.g., toggling switches on and off for specific DERs), and sometimes continuous spaces (e.g., adjusting setpoints for selected DERs) [57, 54].
RL has been applied to optimize restoration strategies, such as scheduling repairs for components [60], dispatching mobile energy resources [61, 62], and managing repair crews [62]. [60] compares the performance of the Q-learning algorithm with Greedy Search (GS) and Exhaustive Search (ES). Q-learning outperforms GS and ES by achieving optimality faster.
Extreme events can disrupt communication infrastructure, making centralized decision-making impractical. To address this, a decentralized framework is proposed in [62] for deploying DERs. The proposed method outperforms other models in the resilience index by 6-9%. However, it does not account for uncertainties in DERs, and its optimal performance remains 3.53% lower than mixed-integer linear programming (MILP).
Beyond RL, other ML approaches are also instrumental in enhancing power system resilience. Predictive control policies for MG management have employed tree-based models to forecast load demands and solar production [63]. Bayesian Networks and hybrid statistical algorithms like the Fragility-curve Monte Carlo Simulation Scenario Reduction (FC-MC-SCENRED) have been compared for pre-hurricane system optimization for distribution networks [64]. Graph neural networks (GNNs) have been proposed to assess node criticality and identify vulnerable lines in cascading outage studies [65, 66]. For disturbance management, CNNs paired with autoencoders have been developed to dynamically adjust shunt capacitors and transformers during electromagnetic pulse disturbances [67]. A one-dimensional CNN has been designed to detect data intrusions in smart meters [68]. Advanced deep learning techniques have been applied for system-wide online load restoration with high wind power penetration [55], and MG planning in active distribution systems using Bayesian Regularization Backpropagation [69].
Recent advancements in transformer-based generative intelligence offer immense potential to apply in the power systems to better capture complex patterns between parameters. This transformer-based model can tune millions of parameters at a time, which provides the ability to represent more accurate results [96]. The models such as eGridGPT and Weather Foundation Model, are used to train and forecast more accurate power grid analysis and decision making [95]. Domain-specific large language model (LLM) framework extends beyond conventional predictive learning by introducing reasoning, context awareness, and multi-modal understanding of the grid’s characteristics [96]. Transformer-based generative architectures demonstrate that self-attention mechanisms can effectively combine information from different sources and understand how events change over time and locations [99, 98]. However, the model requires high-performance computing infrastructure to process datasets for training and real-time inference. In addition, the model may provide inaccurate output and need verification. Despite these challenges, these models offer a significant opportunity to deliver more accurate and adaptive analyses of power systems behavior under various operating conditions.

### v2

@@ 4 Resilience of Interdependent Power and Community Infrastructure Systems
Modern power systems operate within an interconnected ecosystem of critical infrastructures whose joint performance determines overall resilience. While power system resilience traditionally refers to the grid’s ability to withstand and rapidly recover from HILP events [94], community functionality and broader infrastructure recovery are inherently linked to this capability. Community resilience—defined as the capacity of populations and institutions to prepare for, absorb, and recover from disasters—depends heavily on the continuity of electricity supply and the robustness of interdependent systems such as transportation, water, gas, and telecommunication networks [127]. In highly electrified and infrastructure-reliant societies, disruptions in the power grid can cascade across sectors, amplifying socioeconomic impacts. Accordingly, assessing resilience requires a system-of-systems (SoS) perspective that captures both technical inter-dependencies and community consequences. This perspective highlights that resilience outcomes are not determined solely by the reliability of the power grid itself, but also by how disruptions propagate through interconnected infrastructures and affect communities that depend on these services.
@@ 4.1 Infrastructure Interdependence and System-of-Systems Perspective
The electric grid is tightly coupled with other critical infrastructures (Fig. 4). Under normal conditions, such coupling enhances system efficiency through coordinated control, information sharing, and optimized resource allocation. However, during HILP events, these same inter-dependencies accelerate disruption propagation and amplify cascading failures across sectors [84, 72, 78, 21].
@@
From a resilience perspective, these coupled infrastructures represent both stabilizing supports and potential pathways for systemic collapse. Telecommunications enable situational awareness and remote control; gas and water networks supply essential fuels and cooling resources for generation; and transportation systems facilitate repair crew mobilization, logistics, and power system recovery activities [70]. Recent studies highlight this need by quantifying cross-sectoral sensitivity, recovery delays, and performance degradation under multiple hazards. Table 2 summarizes representative studies examining the interdependence among the electric power systems and other communities and critical infrastructures.
@@
Water infrastructure exhibits strong dependence on electricity for pumping and pressure maintenance. Optimization-based resilience modeling in [90] showed that placing DERs near pump stations can increase Water Distribution System (WDS) resilience by 13.35%. Maintaining pump operation during outages is particularly important because even short power interruptions can disrupt water pressure, sanitation, and emergency response capabilities within affected communities. Under more severe disruptions, potable water systems experience drastic degradation when electrical service is lost. Using the Water Network Tool for Resilience (WNTR) with Pressure-Dependent Demand modeling, [65] demonstrated that hurricane-induced power outages reduce potable water system resilience by nearly 50%. Such performance losses illustrate how electrical failures can quickly propagate into essential public health services.
@@
Dynamic optimal energy flow modeling in [100] showed that incorporating gas storage improves gas system resilience by 29.5%. and that adding backup generators increases heating-system resilience by 15.6% following hurricanes, indicating that cross-sector energy resources can provide critical redundancy during extreme events. Multi-hazard Monte Carlo simulations in [88] further show that wind and earthquake events reduce energy resilience by approximately 7–14%. These cascading effects can significantly extend recovery timelines for communities reliant on multiple energy services.
@@
Transportation infrastructure also relies heavily on the power grid for operational continuity. Seismic resilience simulations in [127] showed that increasing earthquake magnitude from $M=6$ to $M=8$ reduces transportation network resilience by nearly 50%. In large-scale disasters, reduced mobility can therefore prolong outages and slow community recovery processes. Complementary robust optimization studies in [115] demonstrate that strategic hardening of critical power lines and placement of DERs significantly reduce disruption costs and preserve the operability of transportation-critical loads such as traffic signals during HILP events.
@@
Linear regression analysis of post-hurricane recovery in [93] quantified strong cross-sector dependencies during restoration. Water, telecommunications, and hospital systems relied on power recovery for 89%, 93%, and 77% of their restoration trajectories, respectively. Delays in power restoration, therefore, tend to propagate across multiple infrastructures, extending the overall duration of community disruption following extreme events.
These studies indicate that disruptions in the electric grid can propagate across multiple sectors, amplifying societal impacts and delaying community recovery. This SoS perspective suggests that resilience planning should prioritize not only electrical infrastructure reliability but also the protection of services that enable essential community functions such as water supply, transportation, healthcare, and communications.
@@ 4.2 Resilience Hubs for Community Infrastructure Support
Grid resilience strategies are most meaningful when they translate into tangible benefits for the communities they serve. Resilience hubs represent this translation, as they provide a comprehensive, community-centered approach to addressing interdependent, multifaceted vulnerabilities. They recognize the interconnected nature of critical community infrastructure systems and the socioeconomic dimensions through which HILP events affect community members. Resilience hubs integrate public health and safety goals with technical considerations for strengthening power, transportation, communications, and essential services.
Resilience hubs are community-oriented facilities, such as schools, community centers, or libraries, that provide everyday public services during normal operating conditions and serve as disaster shelters, resource centers, and infrastructure support during emergencies [13] (Fig. 5).
@@
Resilience hubs reflect the interdependence of community and grid resilience, addressing both immediate individual social vulnerability concerns and technical system-level infrastructure shortfalls [40]. When strategically designed as integrated techno-physical-social system nodes, they can serve as a foundation for community and grid resilience. For example, a resilience hub may focus on supporting medically vulnerable populations, stabilizing public health services, coordinating emergency response or recovery efforts, ensuring access to cooling or clean water, or enhancing grid reliability using DER and adaptive load management.
Resilience hubs are equipped with resilient backup power systems that not only ensure uninterrupted on-site access to critical services but can also provide grid-supporting functions [40]. This flexible capacity can stabilize nearby interdependent community infrastructure by supporting water pumping stations, emergency communications, transportation assets, and reliable electric vehicle (EV) charging during outages or evacuations. As a result, resilience hubs can be strategically located either to optimize accessibility for residents [41] or near critical infrastructure during grid disruptions [91].
By forming a resilience hub-centered MG using smart switching and grid reconfiguration techniques described in Section 2, operators can support critical infrastructure and services (such as cooling during extreme heat events) during cascading grid failures. Resilience hubs can also coordinate with grid operators to support broader black start recovery. The RL-based reconfiguration methods and outage prediction models discussed in Section 3 further enhance this capability.
The framework in [41] designs and evaluates energy systems for resilience hubs in a region prone to extreme heat, identifying communities with limited access to air conditioning and heightened vulnerabilities to power outages. The resulting resilience hub design incorporates a DER portfolio sized to maintain safe indoor temperatures during a heatwave for up to four days without relying on the grid. Similarly, the grid-aware optimization in [47] demonstrates how hub-centered MGs can mitigate outage impacts by balancing local resilience benefits with broader system constraints. Outage prediction models can guide optimal hub siting by identifying where outage risk, social vulnerability, and infrastructure interdependence converge, while RL-based frameworks can dynamically dispatch emergency responders, manage EV charging, and allocate backup generation in response to evolving event conditions. These examples illustrate how resilience hubs are not merely a community amenity, but the point at which grid resilience technologies, AI-driven prediction, and community vulnerability considerations converge into tangible support for critical infrastructure.
@@
4.3 Socioeconomic and Behavioral Dimensions
A growing body of research has quantified how socioeconomic and behavioral factors shape outage exposure, duration, and subsequent recovery. A data-driven Social Vulnerability Index (SoVI) framework in [68] combined Principal Component Analysis (PCA), Adaptive Random Forests (ARF), and variance-based sensitivity analysis (VBSA) to identify the dominant societal drivers of outage vulnerability. Projected SoVI values for 2030 indicate a reduction in vulnerability disparities compared to 2022, reflecting long-term adaptation and incremental infrastructure improvement. Similar disparities were documented in [46], where a 0.1 increase in socioeconomic vulnerability was associated with a 6.1% increase in outage duration. Counties with the longest restoration delays (+6 hours above the mean) exhibited very high SoVI values ($\text{SoVI}\approx 0.91$). The statistical relationship between SoVI and outage duration highlights how underlying social conditions influence not only the likelihood of disruption but also the pace of infrastructure recovery. Communities with higher vulnerability frequently face constraints such as limited financial resources, aging infrastructure, reduced access to backup energy systems, and weaker institutional support for restoration efforts. Under these circumstances, similar grid disturbances can produce substantially different outage experiences across communities, with vulnerable populations often facing prolonged service interruptions.
Outage exposure also differs across racial and ethnic lines. A multilevel Bayesian analysis in [87] showed that counties with high racial/ethnic vulnerability $(\text{SoVI}>0.9$) experienced the longest outage durations during extreme weather events, including hurricanes, heat waves, and floods. Similarly, a spatially resolved analysis in [31] found that counties with high SoVI experienced approximately two times more outages lasting $\geq 1$ hour and approximately three times more outages lasting $\geq 8$ hours compared to low-SoVI counties. These patterns illustrate that outage impacts are not determined solely by the intensity of the hazard or the configuration of the power grid. Instead, pre-existing socioeconomic inequalities influence exposure to disruptions, access to preparedness resources, and the ability of communities to respond effectively when outages occur. As a result, communities with higher vulnerability often experience both greater outage frequency and slower restoration timelines.
Behavioral factors further amplify these disparities. A survey study in New York City [32] found that although 58% of respondents reported feeling generally prepared for outages, only 32% felt capable of enduring multi-day interruptions. Users of durable medical equipment (DME) were more prepared (70%), yet 74% still feared serious health complications during outages. These findings reveal a crucial gap between perceived and actual preparedness. Additional household-level evidence from [8] shows that minority households reported significantly longer outages, while lower-income households experienced almost twice as many outage events. Such disparities suggest that the societal burden of power outages is unevenly distributed across populations. Differences in preparedness, resource availability, and household resilience capacity can significantly influence how communities experience and recover from disruptions. Table 2 summarizes the interdependency between power systems resilience and SoVI.
Age and health status further influence outage-related risk. Children, older adults, and medically fragile populations experience disproportionately severe impacts from prolonged outages and associated cascading failures [24, 56]. Documented public health risks include cardiovascular events, respiratory complications, renal failure, heat stress, gastrointestinal illness, and increased mental health burdens as outage durations rise [24]. Because electricity supports critical services such as medical equipment, climate control, and healthcare infrastructure, prolonged outages can quickly translate into severe health risks for vulnerable populations. To mitigate these outcomes, recent work highlights the importance of community-scale renewable energy resources and medically focused resilience planning to maintain essential electricity access during HILP events [56]. At the same time, emerging socio-technical models increasingly incorporate behavioral and human-response data [106]. Across these studies, a reinforcing feedback mechanism emerges between social vulnerability and power system resilience. Communities with higher vulnerability tend to experience longer outages and slower recovery, which in turn can exacerbate health risks, economic hardship, and reduced access to essential services. These impacts may weaken community preparedness and adaptive capacity, increasing susceptibility to future disruptions and reinforcing existing inequalities.
Recognizing this dynamic has important implications for resilience planning. Infrastructure investments should therefore not rely exclusively on engineering-based risk metrics. Instead, resilience strategies should incorporate equity-centered indicators, when prioritizing resilience investments. Integrating these considerations into grid hardening strategies, DER deployment, and resilience hub planning can help ensure that resilience improvements address the communities most vulnerable to outage impacts.